Saturated steam at 200 bar is to be heated at constant pressure to 650°C. Using appropriate deviations, determine ∆h, the heat added, ∆s, and the specific volume at state 2. Compare results with values read from the Mollier chart, Item B 16 (SI).

1
Expert's answer
2021-12-13T11:10:04-0500

Enthalpy

"\u2206H=U+PV"

"∆H=U+200\times10^5\times V"

Entropy

"S=\frac{∆Q}{T}"

"S=\frac{mL}{T}"

boron

"M=50.2\times1.67\times10^{-27}kg"

Latent heat


"S=\frac{50.2\times1.67\times10^{-27}\times10}{923}=9.082\times10^{-28}J/K"

Millier chart